# Theory of local $\mathbb{Z}_{2}$ topological markers for finite and periodic two-dimensional systems

### v1

@@ I Introduction
The topological phases of insulators are characterized by a bulk topological invariant, whose non-trivial character is manifested by gapless boundary modes. In the framework of two-dimensional systems, two main classes of topological insulators (TI) can be identified: the quantum anomalous Hall (also known as Chern) insulators (QAHI) [1] and time-reversal (TR) symmetric TIs, also known as quantum spin Hall insulators (QSHI) [2, 3, 4]. The topological phases of QAHIs are classified according to the Chern number $C\in\mathbb{Z}$, which signals the presence of chiral edge states protected by the non-trivial topology. As for QSHIs, TR symmetry is required and the symmetry-protected topological phase [2] is classified by an index $\nu\in\mathbb{Z}_{2}$. QSHIs display an odd number of helical edge states and edge currents that are spin-momentum locked.
@@
Within periodic boundary conditions (PBC), the presence of a non-zero topological invariant is related to the impossibility of constructing a smooth gauge in the whole Brillouin zone (BZ) [5, 6] and represents a topological obstruction. Topological invariants are usually defined as global quantities of the system, and in general they cannot be evaluated as the expectation value of some operator. The task further complicates when considering symmetry-protected topological phases, as it becomes necessary to include the protecting symmetries in the definition of the topological invariant. Examples for TR-symmetric insulators include the concept of TR polarization [7], the parity of TR invariant momenta [8], or the flow of hybrid Wannier functions centers [6, 9, 10]. PBCs are usually chosen when doing electronic structure simulations of solids, and symmetries are typically directly incorporated in reciprocal space [8, 7], rather than real space. However, by construction, these global formulations target homogeneous and periodic systems. In the case of periodic and disordered systems a number of strategies have been developed to compute the global topological invariant, such as the single-point sampling of the BZ in the large supercell framework [11, 12], the Bott index [13, 14], the structural spillage [15], and methods based on the scattering matrix [16, 17] or on the non-commutative index theorem [18, 19, 20].
@@
When dealing with finite or inhomogeneous systems, global invariants are of no avail, and one must adopt a space-resolved approach, as done for instance in Refs. [21, 22, 23, 24, 25, 26]. The ability to sample these quantities locally in real space can be particularly useful when dealing, for instance, with amorphous systems [27]. The long-range order—that usually allows to calculate the topological invariant via reciprocal space formulations—is lost, and only on the local scale the system appears to retain an ordered arrangement. Since the topological invariant is related to the organization of electrons in the ground-state, in principle it does not require the existence of long-range order or translational invariance. Being the ground-state electronic structure “nearsighted” [28], it should be possible to probe the local topology of a region just with the knowledge on its neighborhood in real space. A simple and relevant scenario is a heterostructure composed by two regions with different topology, such as an insulator in a topological phase at the interface with a trivial insulator (or vacuum). There, the meaning of a global invariant is ambiguous, since a single invariant cannot capture the individual bulk topology of the two regions and the presence of metallic edge states at the interface between them.
@@
In this work, we develop space-resolved formulas for the $\mathbb{Z}_{2}$ topological invariant that can be used to probe the local topology of inhomogeneous and non-crystalline two-dimensional systems. Over the past years, a number of local markers have been developed to investigate the presence of topological phases in different systems. Examples include the local Chern marker [21], that has been used to study also QSHIs [29], the mirror-Chern marker [22], a spin-Chern marker based on the spin-Berry curvature [30], a definition of the local $\mathbb{Z}_{2}$ invariant based on a flux-insertion-induced spectral flow [31], and a local marker to identify crystalline topological phases [32]. In particular, the archetypal case of a local Chern marker for two-dimensional QAHIs has been investigated in Ref. [21] within open boundary conditions (OBCs) and in Ref. [33] within PBCs. Moreover, a generalization of the concept of local marker in odd dimensions $D$ has been introduced in Ref. [34] for both $\mathbb{Z}$-phases (local Chiral marker) and $\mathbb{Z}_{2}$-phases (local Chern-Simons marker). That is done by introducing a $(D+1)$-dimensional local Chern marker, extending the one by Bianco-Resta [21] to higher dimensions, together with a family of parametrized projectors (where the parameter acts as the additional dimension) that have to be integrated in order to get a local marker when $D$ is odd. In particular, the invariant is $\mathbb{Z}$-valued if the single-particle density matrix obeys a chiral constraint, and $\mathbb{Z}_{2}$ otherwise [35]. Recently, a universal topological marker, valid in any dimension and symmetry class, has been proposed in Ref. [36]. This marker has been successfully tested on various systems belonging to different symmetry classes, but in the most general case (of interest for this work) of two-dimensional class AII systems with Rashba spin-orbit coupling (SOC), such as the Kane-Mele model [3], a general and practical formulation of the local $\mathbb{Z}_{2}$ marker that can always be applied is still missing. Indeed, Rashba SOC (that can be intrinsic or induced, for instance, by an electric field perpendicular to the system) violates $S_{z}$ conservation, so one can no longer consider spins to be “up” or “down”, and must resort to a different strategy to calculate the $\mathbb{Z}_{2}$ invariant.
@@
There are many available methods to compute the global topological invariant of QSHIs by making explicit use of TR symmetry [8, 7]. For our purposes, two particularly noteworthy approaches are the spin-Chern number [37], and the calculation of hybrid Wannier charge centers [38, 10]. The underlying common strategy between these two methods is to split the Hilbert space spanned by the occupied states as a sum of subspaces such that the individual Chern number associated to each subspace is an integer [38, 10]. That can be done in an arbitrary way, but the individual Chern numbers are meaningful only when such splitting is realized according to the symmetries protecting the topological phase. In QSHIs, for instance, one can use the projected spin operator [37] or the TR symmetry [39] to obtain two subspaces where TR must be broken in order to allow a non-zero Chern number. Hence, calculating the $\mathbb{Z}_{2}$ topological invariant translates to the problem of identifying a suitable partitioning of the occupied Hilbert space where it is possible to compute integer individual Chern numbers. However, performing such partitioning might be rather non-trivial in practice. Indeed, it has been discussed [40, 37, 41] that an exponentially localized projector commuting with lattice translations is a sufficient condition to yield integer individual Chern numbers that are invariant with respect to small perturbations of the system.
@@
Here, we propose two formulations for local $\mathbb{Z}_{2}$ invariants: the first leverages the spin-Chern number introduced by Prodan [37], while the second is based solely on TR symmetry, in the spirit of the work by Soluyanov and Vanderbilt [39]. These two strategies are used to define markers for finite samples in OBCs as well as for large supercells in PBCs (where the BZ can be effectively sampled by a single point, typically $\Gamma$). Our approach is particularly useful when dealing with non-homogeneous systems, such as disordered samples and heterostructures, giving the possibility of inspecting the topology locally in real space.
@@
The paper is organized as follows. In Sec. II we recap on some relevant results from the literature, namely the OBC and PBC local Chern markers developed for QAHIs. In Sec. III, we first discuss the local spin-Chern marker and then introduce the local $\mathbb{Z}_{2}$ marker based on TR symmetry within OBCs, that we further generalize to PBCs in Sec. IV. Then, in Sec. V we discuss the tight-binding models used to benchmark our markers and introduce a generalized smearing, which is particularly useful when dealing with heterostructures. In Sec. VI, we present numerical results and compare the different methods introduced. Finally, we summarize our results and conclusions in Sec. VII.
@@ II Local Chern markers
In this section we review the OBC and PBC formulations of the LCM introduced in Refs. [21, 33]. Consider a QAHI with $N_{occ}$ occupied bands, and let |$\ket{u_{n\mathbf{k}}}=e^{-i\mathbf{k}\cdot\mathbf{r}}\ket{\psi_{n\mathbf{k}}}$ be the periodic part of Bloch functions. The Hamiltonian of a QAHI breaks TR symmetry, and its topological invariant is the Chern number:
@@
such that its macroscopic average describes the local topology. In a pristine system, that reduces to the trace per unit area $\mathrm{Tr_{A}}$ in the region of interest, and results in the local Chern number of the system. In Equations 2 and 3, $x$ and $y$ are the Cartesian components of the position operator $\mathbf{r}$, and $\mathcal{P}=\sum_{n=1}^{N_{occ}}\ket{u_{n}}\bra{u_{n}}$ | is the ground-state projector. The key feature of this formulation is that, by being expressed as a trace, the Chern number can be evaluated locally in real space, making it particularly suited to study disordered and inhomogeneous systems. The LCM defined in Eq. 3 is also related to the geometrical intrinsic contribution to the local anomalous Hall conductivity in metals [42, 43]. So defined, the LCM of Eq. 3 could, in principle, be applied in both PBCs and OBCs, since the operators $[r_{\alpha},\mathcal{P}]$ are well-defined in both cases, despite $\mathbf{r}$ being an illegitimate operator within PBCs [44]. However, when dealing with numerical simulations, the operators $[r_{\alpha},\mathcal{P}]$ are not directly accessible and the usual way to construct them is to multiply the “standard” position operator $\mathbf{r}$ with $\mathcal{P}$, thus limiting this method to the OBC case only [33]. Equation 3 shows very clearly that a finite system is always globally trivial in a certain sense, as its Chern number is always zero [21] being the trace over the whole sample of commutators of matrices with finite dimensions.
@@
where $S_{mn}(\mathbf{b}_{j})=\braket{u_{m\Gamma}}{e^{-i\mathbf{b}_{j}\cdot\mathbf{r}}}{u_{n\Gamma}}$ is the overlap matrix between the states at $\Gamma$ and the ones at $\mathbf{b}_{j}$ once the periodic gauge is imposed. The dual vectors satisfy $\braket{\tilde{u}_{n\mathbf{b}_{j}}}{u_{m\Gamma}}=\delta_{nm}$ and represent, in the limit of a large supercell, the vectors obtained from a parallel transport procedure from $\Gamma$ to $\mathbf{b}_{j}$ [45, 12]. The single-point formulation is based on a discretization of the covariant derivative, and depending on whether this is approximated by forward or symmetric finite differences, the asymmetric formula [Eq. 4] or the symmetric one [Eq. 6] can be obtained, respectively:
@@
where $\mathcal{P}_{\mathbf{b}_{j}}=\sum_{n=1}^{N_{occ}}\ket{\tilde{u}_{n\mathbf{b}_{j}}}\bra{\tilde{u}_{n\mathbf{b}_{j}}}$ | and $\mathcal{P}_{\Gamma}=\sum_{n=1}^{N_{occ}}\ket{u_{n\Gamma}}\bra{u_{n\Gamma}}$ | are ground-state projectors. In particular, Eq. 7 derives from the asymmetric single-point formulation [Eq. 4], while Eq. 8 derives from the symmetric one [Eq. 6]. As discussed for the OBC case, also in PBCs the local topology is described by the macroscopic average of the LCM. A key feature of this formulation is that the position operator appears only through the exponential $e^{-i\mathbf{b}_{j}\cdot\mathbf{r}}$, thus resulting in a legitimate operator also within PBCs. The PBC LCM offers a simple picture of the topological obstruction in non-crystalline systems: In the trivial ($C=0$) case, one can always have $[\mathcal{P}_{\mathbf{b}_{1}},\mathcal{P}_{\mathbf{b}_{2}}]=0$ due to the existence of a smooth gauge in the whole BZ. In the topological case, however, this construction is impossible, and a non-zero (local) Chern number arises from the fact that $\mathcal{P}_{\mathbf{b}_{1}}$ and $\mathcal{P}_{\mathbf{b}_{2}}$ (locally) do not commute, also in inhomogeneous systems. If one considers the case of a finite sample inside a much larger supercell with PBCs, the LCM of Eq. 7 can be expanded in powers of the linear dimension of the supercell $L$: In the limit of $L\rightarrow\infty$, the expansion converges to the LCM of Eq. 3. Both the OBC and PBC formulations of the LCM of a QAHI can be applied to study the local topology of disordered and amorphous materials, as well as inhomogeneous systems such as trivial/topological heterojunctions and superlattices.
@@ III Local $\mathbb{Z}_{2}$ markers for finite systems
Because of Kramers’ theorem, each eigenvalue of a TR symmetric Hamiltonian is at least doubly degenerate. Hence, only half of the occupied states is really needed to compute the topological invariant, since the other half can be obtained by symmetry. One can then split the occupied Hilbert space into two Chern subspaces and obtain two set of states (built in general as linear combination of the occupied eigenstates of the Hamiltonian) that are mapped onto each other by TR symmetry. For instance, one can split the Kramers-degenerate eigenstates of the Hamiltonian such that, for each vector assigned to a subspace, its TR partner is assigned to the other. By doing so, in each subspace TR symmetry is broken by construction, so it is possible to compute its individual Chern number [38, 10]. Moreover, since TR symmetry forces $C=0$, the two subspaces must be characterized by opposite individual Chern numbers. Here we promote the individual Chern number to a local individual Chern marker, and use Eq. 3 on the states belonging to just one of the two subspaces. Obtaining a decomposition of the occupied Hilbert space that leads to an integer individual Chern number, however, is not a trivial task. In fact, we need exponentially localized projectors onto the subspaces for the individual Chern numbers to be integers [40, 37, 41]. Here we propose two different methods to realize the splitting of the occupied states into two Chern subspaces characterized by exponentially localized projectors, first via the projected spin operator $\mathcal{P}S_{z}\mathcal{P}$ and then by using TR symmetry $\Theta$, leading to two distinct local markers of the $\mathbb{Z}_{2}$ topology.

### v2

@@ I Introduction
The topological phases of insulators are characterized by a bulk topological invariant, whose non-trivial character is manifested by gapless boundary modes. In the framework of two-dimensional systems, two main classes of topological insulators (TI) can be identified: The quantum anomalous Hall (also known as Chern) insulators (QAHI) [1] and the time-reversal (TR) symmetric TIs, also known as quantum spin Hall insulators (QSHI) [2, 3, 4]. The topological phases of QAHIs are classified according to the Chern number $C\in\mathbb{Z}$, which signals the presence of chiral edge states protected by the non-trivial topology. As for QSHIs, TR symmetry is required and the symmetry-protected topological phase [2] is classified by an index $\nu\in\mathbb{Z}_{2}$. QSHIs display an odd number of helical edge states and edge currents that are spin-momentum locked.
@@
Within periodic boundary conditions (PBC), the presence of a non-zero topological invariant is related to the impossibility of constructing a smooth gauge in the whole Brillouin zone (BZ) [5, 6] and represents a topological obstruction. Topological invariants are usually defined as global quantities of the system, and in general they cannot be evaluated as the expectation value of some operator. The task further complicates when considering symmetry-protected topological phases, as it becomes necessary to include the protecting symmetries in the definition of the topological invariant. Examples for TR-symmetric insulators include the concept of TR polarization [7], the parity of TR invariant momenta [8], or the flow of hybrid Wannier functions centers [6, 9, 10]. For electronic structure simulations of solids PBCs are usually chosen, and symmetries are typically directly incorporated in reciprocal space [8, 7], rather than real space. However, by construction, these global formulations target homogeneous and periodic systems. In the case of periodic and disordered systems a number of strategies have been developed to compute the global topological invariant, such as the single-point sampling of the BZ in the large supercell framework [11, 12], the Bott index [13, 14], the structural spillage [15], and methods based on the scattering matrix [16, 17] or on the non-commutative index theorem [18, 19, 20].
@@
When dealing with finite or inhomogeneous systems, global invariants are of no avail, and one must adopt a space-resolved approach, as done for instance in Refs. [21, 22, 23, 24, 25, 26]. The ability to sample these quantities locally in real space can be particularly useful when dealing, for instance, with amorphous systems [27]. The long-range order—that usually allows calculating the topological invariant via reciprocal space formulations—is lost, and only on the local scale the system appears to retain an ordered arrangement. Since the topological invariant is related to the organization of electrons in the ground state, in principle it does not require the existence of long-range order or translational invariance. As the ground-state electronic structure is “nearsighted” [28], it should be possible to probe the local topology of a region just with the knowledge of its neighborhood in real space. A simple and relevant scenario is a heterostructure composed of two regions with different topology, such as an insulator in a topological phase at the interface with a trivial insulator (or vacuum). There, the meaning of a global invariant is ambiguous, since a single invariant cannot capture the individual bulk topology of the two regions and the presence of metallic edge states at the interface between them.
@@
In this work, we develop space-resolved formulas for the $\mathbb{Z}_{2}$ topological invariant that can be used to probe the local topology of inhomogeneous and non-crystalline two-dimensional systems. Over the past years, a number of local markers have been developed to investigate the presence of topological phases in different systems. Examples include the local Chern marker [21], that has been used to study also QSHIs [29], the mirror-Chern marker [22], a spin-Chern marker based on the spin-Berry curvature [30], a definition of the local $\mathbb{Z}_{2}$ invariant based on a flux-insertion-induced spectral flow [31], a local marker to identify crystalline topological phases [32] and a layer-resolved spin-Chern marker for 3D TIs [33]. Recently, a marker to investigate higher-order topological phases has also been proposed [34]. In particular, the archetypal case of a local Chern marker for two-dimensional QAHIs has been investigated in Ref. [21] within open boundary conditions (OBCs) and in Ref. [35] within PBCs. The local Chern marker [21] can also be accessed experimentally by measuring local circular dichroism [36, 37, 38]. Moreover, a generalization of the concept of local marker in odd dimensions $D$ has been introduced in Ref. [39] for both $\mathbb{Z}$-phases (local Chiral marker) and $\mathbb{Z}_{2}$-phases (local Chern-Simons marker). That is done by introducing a $(D+1)$-dimensional local Chern marker, extending the one by Bianco-Resta [21] to higher dimensions, together with a family of parametrized projectors (where the parameter acts as the additional dimension) that have to be integrated in order to get a local marker when $D$ is odd. In particular, the invariant is $\mathbb{Z}$-valued if the single-particle density matrix obeys a chiral constraint, and $\mathbb{Z}_{2}$ otherwise [40]. Recently, a universal topological marker, valid in any dimension and symmetry class, has been proposed in Ref. [41]. This marker has been successfully tested on various systems belonging to different symmetry classes, but in the most general case (of interest for this work) of two-dimensional class AII systems with Rashba spin-orbit coupling (SOC), such as the Kane-Mele model [3], a general and practical formulation of the local $\mathbb{Z}_{2}$ marker that can always be applied is still missing. Indeed, Rashba SOC (that can be intrinsic or induced, for instance, by an electric field perpendicular to the system) violates $S_{z}$ conservation, so one can no longer consider spins to be “up” or “down”, and must resort to a different strategy to calculate the $\mathbb{Z}_{2}$ invariant.
@@
There are many available methods to compute the global topological invariant of QSHIs by making explicit use of TR symmetry [8, 7]. For our purposes, two particularly noteworthy approaches are the spin-Chern number [42] and the calculation of hybrid Wannier charge centers [43, 10]. The underlying common strategy between these two methods is to split the Hilbert space spanned by the occupied states as a sum of subspaces such that the individual Chern number associated to each subspace is an integer [43, 10]. That can be done in an arbitrary way, but the individual Chern numbers are meaningful only when such splitting is realized according to the symmetries protecting the topological phase. In QSHIs, for instance, one can use the projected spin operator [42] or the TR symmetry [44]: The two subspaces obtained individually break TR symmetry and can host a non-zero Chern number. Hence, calculating the $\mathbb{Z}_{2}$ topological invariant translates to the problem of identifying a suitable partitioning of the occupied Hilbert space where it is possible to compute integer individual Chern numbers. However, performing such partitioning might be rather non-trivial in practice. Indeed, it has been discussed [45, 42, 46] that an exponentially localized projector commuting with lattice translations is a sufficient condition to yield integer individual Chern numbers that are invariant with respect to small perturbations of the system.
@@
Here, we propose two formulations for local $\mathbb{Z}_{2}$ invariants: The first leverages the spin-Chern number introduced by Prodan [42], while the second is based solely on TR symmetry, in the spirit of the work by Soluyanov and Vanderbilt [47, 44]. These two strategies are used to define markers for finite samples in OBCs as well as for large supercells in PBCs (where the BZ can be effectively sampled by a single point, typically $\Gamma$). Our approach is particularly useful when dealing with non-homogeneous systems, such as disordered samples and heterostructures, giving the possibility of inspecting the topology locally in real space.
@@
The paper is organized as follows. In Sec. II, we recap on some relevant results from the literature, namely the OBC and PBC local Chern markers developed for QAHIs. In Sec. III, we first discuss the local spin-Chern marker and then introduce the local $\mathbb{Z}_{2}$ marker based on TR symmetry within OBCs, that we further generalize to PBCs in Sec. IV. Then, in Sec. V we discuss the tight-binding models used to benchmark our markers and introduce a generalized smearing, which is particularly useful when dealing with heterostructures. In Sec. VI, we present numerical results and compare the different methods introduced. Finally, we summarize our results and conclusions in Sec. VII.
@@ II Local Chern markers
In this section we review the OBC and PBC formulations of the LCM introduced in Refs. [21, 35]. Let us consider a QAHI with $N_{occ}$ occupied bands, and let |$\ket{u_{n\mathbf{k}}}=e^{-i\mathbf{k}\cdot\mathbf{r}}\ket{\psi_{n\mathbf{k}}}$ be the periodic part of Bloch functions. The Hamiltonian of a QAHI breaks TR symmetry, and its topological invariant is the Chern number:
@@
such that its macroscopic average over a small region in real space describes the local topology. In a pristine system, that reduces to the trace per unit area $\mathrm{Tr_{A}}$ in the region of interest, and results in the local Chern number of the system. In Equations 2 and 3, $x$ and $y$ are the Cartesian components of the position operator $\mathbf{r}$, and $\mathcal{P}=\sum_{n=1}^{N_{occ}}\ket{u_{n}}\bra{u_{n}}$ | is the ground-state projector. The key feature of this formulation is that, by being expressed as a trace, the Chern number can be evaluated locally in real space, that makes it particularly suited to study disordered and inhomogeneous systems. The LCM defined in Eq. 3 is also related to the geometrical intrinsic contribution to the local anomalous Hall conductivity in metals [48, 49]. So defined, the LCM of Eq. 3 could, in principle, be applied in both PBCs and OBCs, since the operators $[r_{\alpha},\mathcal{P}]$ are well-defined in both cases, despite $\mathbf{r}$ being an illegitimate operator within PBCs [50]. However, when dealing with numerical simulations, the operators $[r_{\alpha},\mathcal{P}]$ are not directly accessible and the usual way to construct them is to multiply the “standard” position operator $\mathbf{r}$ with $\mathcal{P}$, thus limiting this method to the OBC case only [35]. Equation 3 shows very clearly that a finite system is always globally trivial in a certain sense, as its Chern number is always zero [21] being the trace over the whole sample of commutators of matrices with finite dimensions.
@@
where $S_{mn}(\mathbf{b}_{j})=\braket{u_{m\Gamma}}{e^{-i\mathbf{b}_{j}\cdot\mathbf{r}}}{u_{n\Gamma}}$ is the overlap matrix between the states at $\Gamma$ and the ones at $\mathbf{b}_{j}$ once the periodic gauge is imposed. The dual vectors satisfy $\braket{\tilde{u}_{n\mathbf{b}_{j}}}{u_{m\Gamma}}=\delta_{nm}$ and represent, in the limit of a large supercell, the vectors obtained from a parallel-transport procedure from $\Gamma$ to $\mathbf{b}_{j}$ [51, 12]. The single-point formulation is based on a discretization of the covariant derivative, and depending on whether this is approximated by forward or symmetric finite differences, the asymmetric formula [Eq. 4] or the symmetric one [Eq. 6] can be obtained, respectively:
@@
where $\mathcal{P}_{\mathbf{b}_{j}}=\sum_{n=1}^{N_{occ}}\ket{\tilde{u}_{n\mathbf{b}_{j}}}\bra{\tilde{u}_{n\mathbf{b}_{j}}}$ | and $\mathcal{P}_{\Gamma}=\sum_{n=1}^{N_{occ}}\ket{u_{n\Gamma}}\bra{u_{n\Gamma}}$ | are ground-state projectors. In particular, Eq. 7 derives from the asymmetric single-point formulation [Eq. 4], while Eq. 8 derives from the symmetric one [Eq. 6]. As discussed for the OBC case, also in PBCs the local topology is described by the macroscopic average of the LCM over a small region in real space. A key feature of this formulation is that the position operator appears only through the exponential $e^{-i\mathbf{b}_{j}\cdot\mathbf{r}}$, thus resulting in a legitimate operator also within PBCs. The PBC LCM offers a simple picture of the topological obstruction in non-crystalline systems: In the trivial ($C=0$) case, one can always have $[\mathcal{P}_{\mathbf{b}_{1}},\mathcal{P}_{\mathbf{b}_{2}}]=0$ due to the existence of a smooth gauge in the whole BZ. In the topological case, however, this construction is impossible, and a non-zero (local) Chern number arises from the fact that $\mathcal{P}_{\mathbf{b}_{1}}$ and $\mathcal{P}_{\mathbf{b}_{2}}$ (locally) do not commute, also in inhomogeneous systems. If one considers the case of a finite sample inside a much larger supercell with PBCs, the LCM of Eq. 7 can be expanded in powers of the linear dimension of the supercell $L$: In the limit of $L\rightarrow\infty$, the expansion converges to the LCM of Eq. 3. In addition, while the OBC LCM [Eq. 3] vanishes upon tracing over the entire system, the PBC LCM [Eqs. 7,8] results in the single-point Chern number [11], even in the disordered case. Both the OBC and PBC formulations of the LCM can be applied to study the local topology of disordered and amorphous materials, as well as inhomogeneous systems such as trivial/topological heterojunctions and superlattices.
@@ III Local $\mathbb{Z}_{2}$ markers for finite systems
Because of Kramers’ theorem, each eigenvalue of a TR symmetric Hamiltonian is at least doubly degenerate. Hence, only half of the occupied states is really needed to compute the topological invariant, since the other half can be obtained by symmetry. One can then split the occupied Hilbert space into two Chern subspaces and obtain two set of states (built in general as linear combination of the occupied eigenstates of the Hamiltonian) that are mapped onto each other by TR symmetry. For instance, one can split the Kramers-degenerate eigenstates of the Hamiltonian such that, for each vector assigned to a subspace, its TR partner is assigned to the other. By doing so, in each subspace TR symmetry is broken by construction, so it is possible and meaningful to compute its individual Chern number [43, 10]. Moreover, since TR symmetry forces $C=0$, the two subspaces must be characterized by opposite individual Chern numbers. Here we promote the individual Chern number to a local individual Chern marker, and use Eq. 3 on the states belonging to just one of the two subspaces. Obtaining a decomposition of the occupied Hilbert space that leads to an integer individual Chern number, however, is not a trivial task. In fact, for the individual Chern numbers to be well-defined, we need exponentially localized projectors onto their corresponding subspaces  [45, 42, 46]. Here, we propose two different methods to realize the splitting of the occupied states into two Chern subspaces characterized by exponentially localized projectors, first via the projected spin operator $\mathcal{P}S_{z}\mathcal{P}$ and then by using the TR symmetry operator $\Theta$, leading to two distinct local markers of the $\mathbb{Z}_{2}$ topology.